OpenStudy (anonymous):

x^2+y^2-4x+4y-17=0 find the vertex and radius of the circle

OpenStudy (anonymous):

To do this properly you have to rewrite it in the form: \[(x-h)^2 + (y-k)^2 = r^2\]

myininaya (myininaya):

group your x terms together group your y terms together complete the square for your x's x^2-4x+(-4/2)^2 complete the square for your y's y^2+4y+(4/2)^2 see how i took b and divided it by 2 and then squared find c so that it completes the square g^2+bg+c g^2+bg+(b/2)^2=(g+b/2)^2 this should help let me know if you run into any problems